Question 12

Use the interactive website to determine the percent of cases that fall below a raw score of 15 in the distribution described in Question 9 (Mean = 50 SD = 15).

a. .98%

b. 1.86%

c. 3.50%

d. 4.77%

Question 13

Use the interactive website to determine the percentage of cases that fall in between the raw scores of 15 and 30 for the distribution described in Question 9 (Mean = 50 SD = 15).

a. 7.18%

b. 7.58%

c. 7.77%

d. 8.14%

Question 14

In the variable distribution described in Question 9 (Mean = 50 SD = 15), what raw score corresponds to a z score of -2.5?

a. 7.5

b. 10

c. 12.5

d. 15

Question 15

In the variable distribution described in Question 9 (Mean = 50 SD = 15), what would be the raw score that corresponds to the z score of 0?

a. 0

b. 15

c. 30

d. 50

Question 16

You draw a random sample of 626 Fort Hays State students and measure their Grade Point Average. The mean GPA of the sample is 3.0, with a standard deviation of .5. What is the standard error of the mean for this sampling distribution of means?

a. .01

b. .02

c. .03

d. .04

Question 17.

You draw a random sample of 626 Fort Hays State students and measure their Grade Point Average. The mean GPA of the sample is 3.0, with a standard deviation of .5. What would be the 95% confidence interval for the GPA of the population of all Fort Hays State students.

a. 2.759 to 3.241

b. 2.815 to 3.185

c. 2.881 to 3.119

d. 2.96 to 3.04

Question 18

Question 18

    You draw a random sample of 626 Fort Hays State students and measure their Grade Point Average. The mean GPA of the sample is 3.0, with a standard deviation of .5. What would be the 99% confidence interval for the GPA of the population of all Fort Hays State students.

a. 2.95 to 3.05

b. 2.81 to 3.18

c. 2.74 to 3.26

d. 2.45 to 3.55

Question 19

A random sample of 226 houses in Hays were appraised. The mean value of the sample of houses was $83,510, with a standard deviation of $26,625. What would be standard error of the mean for this sampling distribution of means?

a. $1,350

b. $1,500

c. $1,775

d. $1,900

question 20

A random sample of 226 houses in Hays were appraised. The mean value of the sample of houses was $83,510, with a standard deviation of $26,625. What would be the 95% confidence interval for the mean value of all houses in Hays?

a. $82,900 to $84,730

b. $81,790 to $85,920

c. $80,031 to $86,989

d. $79,550 to 87,470

Question 21

A random sample of 226 houses in Hays were appraised. The mean value of the sample of houses was $83,510, with a standard deviation of $26,625. What would be the 99% confidence interval for the mean value of all houses in Hays?

a. $ $77,005 to $86,115

b. $78,655 to $86,665

c. $79,305 to $87,015

d. $78,948 to $88,072
Answer #1

12)

Mean = 50, SD = 15

P(X < 15) = P{Z < (15 - 50)/15} = P(Z < -2.333) = 0.0098 = 0.98%

13)

P(15 < X < 30) = P(-2.333 < Z < -1.333) = 8.14%

14)

Raw score = 50 - 2.5*15 = 12.5

15)

Raw score = 50

16)

Standard error = 0.5/√626 = 0.02

17)

The 95% confidence interval is

(3 - 1.96*0.02, 3 + 1.96*0.02) = (2.96, 3.04)

18)

The 99% confidence interval is

(3 + 2.575*0.02, 3 + 2.575*0.02) = (2.95, 3.05)

19)

Standard error = 26,625/√226 = 1771 ≈ $1,775

20)

The 95% confidence interval is

(83,510 - 1.96*1771, 83510 + 1.96*1771)

= $80,031 to $86,989

21)

The 99% confidence interval is

(83,510 - 2.575*1771, 83510 + 2.575*1771)

= $78,948 to $88,072